OpenStudy (anonymous):

Find the equation of the parabola determined by the condition: Axis horizontal, focus (0,0), latus rectum 8. Would someone kindly help me please? :)) Would gladly appreciated.

OpenStudy (anonymous):

so a=4

OpenStudy (anonymous):

focal length is 4

OpenStudy (anonymous):

the vertex is ( 4,0 )

OpenStudy (anonymous):

(y-k)^2 = -4a(x-h)

OpenStudy (anonymous):

y^2 = - 16(x-4)

OpenStudy (anonymous):

if the latus rectum is x=8 , then the distance from the latus rectum to the focus is twice the focal length

OpenStudy (anonymous):

I see... :)

OpenStudy (anonymous):

also, since the focus is always "inside " the parabola, you know the parabola opens to the left

OpenStudy (anonymous):

noted.